### v1

@@ 3.2.2 Soft Prompt
For triple classification task, Given $v$ and $w$, predict $e$,
[…]
For entity prediction task, depending on the missing entity, prompts are created to predict either $v$ or $w$, tbe prompt for predicting tail entity $w$ is:
[…]
The prompt for predicting head entity $v$ is:
[…]
Finally, uses instruction tuning with prompts and expected responses to fine-tune the models. This aligns the predictions with the factual correctness required for KG tasks.
KGT5 52 is pre-trained and fine-tuned on Knowledge Graph Completion (KGC) and Knowledge Graph Question Answering (KGQA) tasks. The model employs regularization to maintain link prediction accuracy. The advantage of this method is its flexibility in handling both KGC and KGQA, making it a versatile solution. However, the complexity of training and fine-tuning a Seq2Seq model can be computationally expensive and time-consuming.
KG-S2S 51 adopts a Seq2Seq generative framework to manage various types of KGC tasks by translating graph structures into flat text representations. This approach allows for unified processing but might suffer from a loss of structural information, which is critical for accurate graph completion.
CSPromp-KG 32 leverages Conditional Soft Prompts to guide pre-trained language models in balancing the integration of structural and textual information for KGC. The method enhances model performance by ensuring a more nuanced understanding of graph structures. The major drawback is the potential overfitting to specific knowledge graph structures, which may limit the generalizability.
The strength of Seq2Seq models lies in their flexibility and effectiveness in generating textual output that adheres to the structural rules of knowledge graphs. However, these models can be prone to generating implausible links if not properly constrained or if the training data is not representative enough of the test scenarios.
3.3 Decoder-Based Methods
In these studies, LLMs are adapted to enhance reasoning and inference capabilities over knowledge graphs, particularly for complex question answering and fact verification tasks. Techniques include adapting attention mechanisms to focus on relevant subgraphs and employing reasoning strategies that leverage structured graph data.
3.3.1 Multi-Level Encoder
Fig. LABEL:fig2.2 illustrates the workflow of this approach.
This category includes models that enhance the capabilities of LLMs in complex reasoning and question answering tasks over knowledge graphs. Techniques involve structuring the input to LLMs in a way that mimics reasoning or using retrieval-augmented generation to enhance the context available for decision-making.
The approach starts by converting a relevant subgraph $\mathcal{G}_{q}$ from the knowledge graph into a sequence using a breadth-first search (BFS)-based serialization. This serialized form is then processed by the PLM. The serialized subgraph $\mathcal{G}_{q}$ is denoted as:
$\mathcal{SG}_{q}=\{v,e,w,\ldots\}$ (21)
where $v$, $e$, and $w$ are nodes and edges in the subgraph.
As for subgraph-aware self-attention, the mechanism is designed to mimic the processing of a GNN, enabling the model to understand the structure of the subgraph and reason over it. It utilizes attention mechanisms to model the relations and interactions between different entities (nodes) and relations (edges) within the subgraph. The attention process can be represented as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}+M\right)V$ (22)
Here, $Q$, $K$, and $V$ are query, key, and value matrices, respectively, and $M$ is a mask matrix that directs attention flow to maintain structural integrity.
The PLM is adapted to handle the serialized subgraph and question data simultaneously, facilitating deep interaction and knowledge sharing between the question context and the graph structure. This integration enables the PLM to perform reasoning tasks directly on the subgraph structure, harnessing both the rich semantic understanding of the PLM and the structured reasoning capabilities typically associated with GNNs.
The PLM undergoes adaptation tuning using a dataset created from synthesized questions and subgraphs. This adaptation helps the PLM better understand and process the unique format of KG data. The fine-tuning phase optimizes the PLM on downstream question answering tasks over KGs, improving accuracy and efficiency in real-world applications.
ReSKGC 81 introduces a retrieval-enhanced Seq2Seq model that uses a retrieval module to enhance the generation of missing entities by fetching semantically relevant triplets. This method improves accuracy and relevance but relies heavily on the effectiveness of the retrieval process, which can be a bottleneck if not optimized properly.
StructGPT 88 employs an Iterative Reading-then-Reasoning framework to enhance LLMs’ reasoning capabilities over structured data like knowledge graphs. It significantly improves reasoning performance, but the iterative process can be resource-intensive and slow, particularly with complex queries.
These models excel in tasks requiring deep reasoning and can handle nuanced queries effectively by leveraging both the structured information from knowledge graphs and the unstructured textual data processed by LLMs. The main drawback is their potential computational inefficiency and the complexity of integrating structured reasoning with generative language models.
3.3.2 Prompt Engineering
Research in this area focuses on integrating comprehensive contextual information and structured knowledge into the training and operation of LLMs. This enables the models to better understand and generate responses based on both the textual descriptions and the underlying graph structures of the knowledge graph. The conventional method flowchart can be seen in Fig. 9.
Figure 9: An overview of prompt engineering method.
Methods under this heading focus on integrating extensive contextual information and structured knowledge graph data within the processing capabilities of LLMs. This often involves using knowledge prompts or other forms of guided input to enhance the LLM’s understanding of the graph structure and semantics.
Complex questions are decomposed into simpler sub-questions based on predefined templates. This is achieved through a slot-filling mechanism driven by LLMs, formalized as:
$\{\mathcal{S}_{i,t}\}_{t=1}^{T}=\mathrm{LLM}(\mathcal{Q}_{i})$ (23)
where $\mathcal{Q}_{i}$ is the initial complex question, and $\mathcal{S}_{i,t}$ represents the decomposed sub-questions.
For each decomposed sub-question, the framework retrieves relevant entities and facts from the KG using predefined logical chains that correspond to the sub-questions, the retrieval process is summarized as:
$\mathcal{C}_{t}=\bigcup_{l=1}^{L_{i}}\{(s,r,o)_{l}\mid(s,r,o)_{l}\in\mathcal{G}\}$ (24)
where $(s,r,o)_{l}$ are the triples retrieved based on the logical chains, with $s$, $r$, and $o$ denoting the subject, relation, and object, respectively.
The candidate reasoning stage involves selecting the correct answers from the retrieved candidates by evaluating their relevance to the sub-questions, expressed using the following function:
$\mathcal{A}_{i}=\mathrm{LLM}(\mathcal{Q}_{i,t}\mid\mathcal{C}_{t})$ (25)
where $\mathcal{A}_{i}$ are the answers derived for sub-question $\mathcal{Q}_{i,t}$, based on candidates $\mathcal{C}_{t}$.
Finally, the responses from all sub-questions are aggregated to construct a comprehensive answer to the original question, which formalized as:
$\text{Response}=\text{Generate}(\{\mathcal{A}_{i}\}_{i=1}^{N})$ (26)
where the function Generate compiles the individual answers into a final coherent response.
KoPA 56 introduces the Knowledge Prefix Adapter (KoPA), which enhances the integration of structural KG information into LLMs for better structural-aware reasoning. While it improves reasoning accuracy, the adaptation of LLMs to specific knowledge graph structures could limit the model versatility across different KGs.
CP-KGC 31 explores the use of zero-shot large language models to enhance text-based KGC without fine-tuning. The advantage is the minimal training requirement, but the reliance on zero-shot capabilities may result in less precise or contextually inappropriate text generation.
KICGPT 53 integrates a triple-based knowledge graph completion retriever with a large language model to address information scarcity in long-tail entities. This method effectively utilizes in-context learning strategies, but its success heavily depends on the quality and relevance of the retrieved knowledge.
These approaches allow LLMs to leverage their pre-trained knowledge effectively by guiding them with structured inputs, enhancing their performance on knowledge-intensive tasks. However, the success of these methods heavily depends on the quality and the extent of the integration between text and graph data, which can be challenging to optimize.
@@ 4 Experiments and Evaluations
In this section, we present a comprehensive suite of experiments and evaluations to assess the performance and effectiveness of the various LLM-enhanced KRL methods discussed previously. We utilize a variety of datasets and metrics to provide a thorough analysis, ensuring that our findings are robust and generalizable. The section is structured to cover the datasets used, the evaluation metrics employed, and the downstream tasks for which the models were tested.
@@ 4.1 Datasets
Dataset # Entity # Relation # Train # Dev # Test Task Year Source FIGER - - 2,000,000 10,000 563 ET 2015 1 Open Entity - - 2,000 2,000 2,000 ET 2018 2 FewRel 1.0 - 80 8,000 16,000 16,000 RC 2018 3 TACRED - 42 68,124 22,631 15,509 RC 2017 4 WN11 38,696 11 112,581 2,609 10,544 TC 2013 5 WN9 6,555 9 11,741 1,337 1,319 TC 2022 6 FB13 75,043 13 316,232 5,908 23,733 TC 2013 5 FB15K-237N 13,104 93 87,282 7,041 8,226 TC 2022 7 FB15K 14,951 1,345 483,142 50,000 59,071 LP 2013 8 FB15k-237 14,541 237 272,115 17,535 20,466 LP 2015 9 UMLS 135 46 5,216 652 661 LP 2018 10 Nations 14 55 1,592 199 201 LP 2007 11 Diabetes 7,886 67 56,830 1,344 1,936 LP 2022 2 YAGO3-10 103,222 30 490,214 2,295 2,292 LP 2018 10 Wikidata5M (transductive) 4,594,485 822 20,614,279 5,163 5,133 LP 2021 13 Wikidata5M (inductive) 4,594,458 822 20,496,514 6,699 6,894 LP(ZS) CoDEx-S 45,869 68 32,888 1,827 1,828 LP 2020 14 CoDEx-M 11,941 50 185,584 10,310 10,311 LP CoDEx-L 45,869 69 551,193 30,622 30,622 LP NELL-ONE 68,545 822 189,635 1,004 2,158 LP(ZS) 2018 15 WN18RR 40,943 11 86,835 3,034 3,134 LP, LP(ZS) 2018 10
@@
  • WN18. WN18 is a benchmark dataset based on the WordNet knowledge graph. It contains a subset of entities and relations, often used for evaluating models on the task of link prediction in knowledge graphs.
@@ 4.3.1 Entity Typing
Model Year FIGER Open Entity Choi et al. Acc Macro Micro P R F1 P R F1 NFGEC (Attentive) 70 2016 54.53 74.76 71.58 - - - - - - NFGEC (LSTM) 70 2016 55.6 75.15 71.73 68.8 53.3 60.1 - - - UFET 98 2018 - - - 77.4 60.6 68 68.8 53.3 60.1 BERT 18 2018 52.04 75.16 71.63 76.37 70.96 73.56 76.4 71 73.6 RoBERTa 83 2019 - - - 77.4 73.6 75.4 - - - ERNIE 38 2019 57.19 76.51 73.39 78.42 72.9 75.56 78.4 72.9 75.6 KnowBert 55 2019 - - - 78.7 72.7 75.6 78.6 73.7 76.1 KEPLER 47 2021 - - - 77.8 74.6 76.2 - - -
@@ 4.3.2 Relation Classification
Model Year FewRel 1.0 TACRED P R F1 P R F1 CNN 23 2015 69.51 69.64 69.35 70.3 54.2 61.2 PA-LSTM 73 2017 - - - 65.7 64.5 65.1 C-GCN 21 2018 - - - 69.9 63.3 66.4 BERT_Base 18 2018 85.05 85.11 84.89 67.23 64.81 66 BERT_Large 18 2018 - - - - - 70.1 RoBERTa 83 2019 - - - 70.4 71.1 70.7 MTB 65 2019 - - - - - 71.5 ERNIE 38 2019 88.49 88.44 88.32 69.97 66.08 67.97 KnowBert 55 2019 - - - 71.6 71.4 71.5 KEPLER 47 2021 - - - 71.5 72.5 72
@@ 4.3.3 Relation Prediction
Model Year FB15K YAGO3-10-100 MR Hits@1 Hits@1 TransE 92 2013 2.5 84.3 - TransR 96 2015 2.1 91.6 - DKRL 36 2016 2 90.8 - TKRL 90 2016 1.7 92.8 - PTransE 76 2015 1.2 93.6 - SSP 87 2017 1.2 - - ProjE 75 2017 1.2 95.7 - KG-BERT 48 2019 1.2 96 - KGT5 52 2022 - - 60 ChatGPT 2023 - - 39 GPT-4 2023 - - 56 LLaMA-7B 59 2023 - - 13 LLaMA-13B 59 2023 - - 1 KG-LLM 50 2023 - - 71
@@ 4.3.4 Triple Classification
Model Year WN11 FB13 UMLS CoDEX-S FB15k-237N Acc Acc Acc Acc Acc NTN 71 2013 86.2 90 - - - TransE 92 2013 75.9 81.5 78.1 72.1 - DistMult 34 2014 87.1 86.2 86.8 66.8 - ComplEx 25 2016 - - 90.8 67.8 65.7 RotatE 84 2019 - - 92.1 75.7 68.5 KG-BERT 48 2019 93.5 90.4 89.7 77.3 56 KG-GPT2 49 2021 85 89 - - - LaSS 57 2021 94.5 91.8 - - - KGT5 52 2022 72.8 66.3 - - - LMKE 60 2022 - 91.7 92.4 - - GPT-3.5 2023 - - 67.6 54.7 60.2 LLaMA-7B 59 2023 21.1 9.1 - - - LLaMA-13B 59 2023 28.1 17.6 - - - KG-LLM 50 2023 95.6 90.2 86 80.3 80.5 KoPA 56 2023 - - 92.6 82.7 77.7
@@ 4.3.5 Link Prediction
Table 9: Experimental statistics of WN18RR and FB15k-237 datasets under the LP subtask in transductive settings.
Model Year WN18RR FB15k-237 MR MRR Hits@1 Hits@3 Hits@10 MR MRR Hits@1 Hits@3 Hits@10 TransE 92 2013 2300 24.3 4.3 44.1 53.2 223 27.9 19.8 37.6 47.4 TransH 94 2014 2524 - - - 50.3 255 - - - 48.6 DistMult 34 2014 3704 44.4 41.2 47 50.4 411 28.1 199 30.1 44.6 TransR 96 2015 3166 - - - 50.7 237 - - - 51.1 TransD 91 2015 2768 - - - 50.7 246 - - - 48.4 ComplEx 25 2016 3921 44.9 40.9 46.9 53 508 27.8 19.4 29.7 45 ConvE 29 2018 4464 45.6 41.9 47 53.1 245 31.2 22.5 34.1 49.7 ConvKB 30 2018 2554 24.9 - - 52.5 257 24.3 - - 51.7 R-GCN 82 2018 6700 12.3 8 13.7 20.7 600 16,4 10 18.1 41.7 KBGAN 46 2018 - 21.5 - - 48.1 - 27.7 - - 45.8 RotatE 84 2019 3340 47.6 42.8 49.2 57.1 177 33.8 24.1 37.5 53.3 KBAT 45 2019 1921 41.2 - - 55.4 270 15.7 - - 33.1 CapsE 19 2019 718 41.5 - - 55.9 403 15 - - 35.6 QuatE 78 2019 3472 48.1 43.6 50 56.4 176 31.1 22.1 34.2 49.5 TuckER 97 2019 - 47 44.3 48.2 52.6 - 35.8 26.6 39.4 54.4 HAKE 44 2019 - 49,7 45.2 51.6 58.2 - 34.6 25 38.1 54.2 AttH 17 2020 - 48.6 44.3 49.9 57.3 - 34.8 25.2 38.4 54 REFE 79 2020 - - - - 56.1 - - - - 54.1 GAATs 39 2020 1270 - - - 60.4 187 - - - 65 Complex-DURA 26 2020 57.1 - - - - 56 - - - - DensE 33 2020 3052 49.1 44.3 50.8 57.9 169 34.9 25.6 38.4 53.5 LineaRE 58 2020 1644 49.5 45.3 50.9 57.8 155 35.7 26.4 39.1 54.5 RESCAL-DURA 80 2020 - 49.8 45.5 - 57.7 - 36.8 27.6 - 55 CompGCN 24 2020 - 47.9 44.3 49.4 54.6 - 35.5 26.4 39 53.5 NePTuNe 69 2021 - - - - 55.7 - - - - 54.7 ComplEx-N3-RP 27 2021 - - - - 58 - - - - 56.8 ConE 28 2021 - 49.6 45.3 51.5 57.9 - 34.5 24.7 38.1 54 Rot-Pro 85 2021 - 45.7 39.7 48.2 57.7 - 34.4 24.6 38.3 54 QuatDE 77 2021 1977 48.9 43.8 50.9 58.6 90 36.5 26.8 40 56.3 NBFNet 68 2021 - 55.1 49.7 - 66.6 - 41.5 32.1 - 59.9 KG-BERT 48 2019 97 21.6 4.1 30.2 52.4 153 23.7 16.9 26 42.7 MTL-KGC 66 2020 89 33.1 20.3 38.3 59.7 132 26.7 17.2 29.8 45.8 Pretrain-KGE 74 2020 - 48.8 43.7 50.9 58.6 - 35 25 38.4 55.4 StAR 88 2021 51 40.1 24.3 49.1 70.9 1117 29.6 20.5 32.2 48.2 MEM-KGC 63 2021 - 57.2 48.9 62 72.3 - 34.9 26 38.2 52.4 LaSS 57 2021 35 - - - 78.6 108 - - - 53.3 SimKGC 86 2021 - 66.7 58.8 72.1 80.5 - 33.6 24.9 36.2 51.1 LP-BERT 62 2022 92 48.2 34.3 56.3 75.2 154 31 22.3 33.6 49 Nayyeri et al. 67 2022 - - - - - - 33.6 24.5 37 52.2 KGT5 52 2022 - 54.2 50.7 - 60.7 - 34.3 25.2 - 37.7 OpenWorld KGC 72 2022 - 55.7 47.5 60.4 70.4 - 34.6 25.3 38.1 53.1 Lovelace et al. 61 2022 - 59.1 51.8 61.6 73.5 - 37.5 28.3 40.8 56.4 LMKE 60 2022 79 61.9 52.3 67.1 78.9 141 30.6 21.8 33.1 48.4 GenKGC 40 2022 - - 28.7 40.3 53.5 - - 19.2 35.5 43.9 KG-S2S 51 2022 - 57.4 53.1 59.5 66.1 - 33.6 25.7 37.3 49.8 kNN-KGE 54 2023 - 57.9 52.5 - - - 28 37.3 - - CSPromp-KG 32 2023 - 57.5 52.2 59.6 67.8 - 35.8 26.9 39.3 53.8 GPT-3.5 2023 - - 19 - - - - 23.7 - - LLaMA-7B 59 2023 - - 8.5 - - - - - - - LLaMA-13B 59 2023 - - 10 - - - - - - - KG-LLM 50 2023 - - 25.59 - - - - - - - CD 20 2024 - 57.6 52.6 60.7 67.2 - - - - - CP-KGC 31 2024 - 67.3 59.9 72.1 80.4 - 33.8 25.1 36.5 51.6 KICGPT 53 2024 - 56.4 47.8 61.2 67.7 - 41.2 32.7 44.8 55.4

### v2

@@ 3.3.2 Prompt Engineering
The LLM then generates a response that indicates the plausibility of the triple. The model uses a scoring function to quantify the plausibility of the triple:
$s(h,r,t)=\text{LLM}(p(h,r,t))$ (27)
where $p(h,r,t)$ is the prompt generated for the triple $(h,r,t)$, and LLM is the large language model.
The training objective includes a cross-entropy loss to fine-tune the LLM on labeled triples:
$L=-\sum_{(h,r,t)\in D}[y\log s(h,r,t)+(1-y)\log(1-s(h,r,t))]$ (28)
where $D$ is the set of training triples, and $y$ is the label indicating whether the triple is true or false.
The KICGPT [41] model integrates a LLM with KGC methods to address challenges in traditional KGC approaches. It significantly reduces training overhead by utilizing in-context learning strategies without requiring explicit model fine-tuning. The model leverages the LLM’s extensive pre-trained knowledge base, combined with a structure-aware KG retriever, to improve the handling of long-tail entities. This integration allows the KICGPT to effectively utilize both structured KG information and the broad knowledge base of the LLM, providing a more robust framework for KGC tasks. The model is particularly innovative in its use of ‘Knowledge Prompt’ strategies, which guide the LLM through structured prompts that include KG information, enhancing the model’s ability to make informed predictions about missing entities in KG.
3.3.3 Structural Fine-tuning
Structural fine-tuning combines embeddings from both structural and textual information, feeding them into the decoder. This approach aims to optimize the decoder’s output by providing a more comprehensive and nuanced representation of the knowledge graph triples. Fig. 10 depicts a structural fine-tuning model KoPA [42].
Figure 10: An overview of structural fine-tuning method.
This method integrates structural embeddings of entities and relations into LLM to enhance knowledge graph completion tasks.
The core idea of KoPA involves two main steps: pre-training structural embeddings and using a prefix adapter to inject these embeddings into the LLM. The structural embeddings are first learned through a self-supervised pre-training process, capturing the structural information of the knowledge graph. The process is described as follows:
For each triple $(h,r,t)$, the structural embeddings $\mathbf{h},\mathbf{r},\mathbf{t}$ are learned using the scoring function:
$F(h,r,t)=\|\mathbf{h}+\mathbf{r}-\mathbf{t}\|^{2}$ (29)
The pre-training objective involves a margin-based ranking loss with negative sampling:
$L_{\text{pre}}=-\log\sigma(\gamma-F(h,r,t))-\sum_{i=1}^{K}\log\sigma(F(h^{\prime}_{i},r^{\prime}_{i},t^{\prime}_{i})-\gamma)$ (30)
where $(h^{\prime}_{i},r^{\prime}_{i},t^{\prime}_{i})$ are negative samples, $\gamma$ is the margin, and $\sigma$ is the sigmoid function.
Once the structural embeddings are pre-trained, KoPA employs a prefix adapter to map these embeddings into the textual token space of the LLM. The transformation is performed as follows:
$\mathbf{K}=P(\mathbf{h})\oplus P(\mathbf{r})\oplus P(\mathbf{t})$ (31)
where $P$ is a projection layer, and $\oplus$ denotes concatenation. The transformed embeddings, referred to as virtual knowledge tokens, are prepended to the input sequence $S$:
$S_{\text{KoPA}}=\mathbf{K}\oplus I\oplus X$ (32)
where $I$ is the instruction prompt, and $X$ is the triple prompt.
The training objective for fine-tuning the LLM with KoPA involves minimizing the cross-entropy loss:
$L_{\text{KoPA}}=-\frac{1}{|S_{\text{KoPA}}|}\sum_{i=1}^{|S_{\text{KoPA}}|}\log P_{\text{LLM}}(s_{i}|s_{<i})$ (33)
where $s_{i}$ are the tokens in the input sequence $S_{\text{KoPA}}$.
KG-GPT2 [43] is a novel adaptation of the GPT-2 language model for KGC. KG-GPT2 leverages the contextual capabilities of GPT-2 to address the challenge of incomplete knowledge graphs by predicting missing links and relationships within them. This model treats each triple in a knowledge graph as a sentence, allowing the GPT-2’s language understanding capabilities to classify the likelihood of these triples accurately. By contextualizing triples in this manner, KG-GPT2 goes beyond traditional embedding techniques, incorporating richer linguistic and semantic information to enhance the graph’s completion.
@@ 4 Experiments and Evaluations
In this section, we present a comprehensive suite of experiments and evaluations to assess the performance and effectiveness of the various LLM enhanced KRL methods discussed previously. We utilize a variety of datasets and metrics to provide a thorough analysis, ensuring that our findings are robust and generalizable. The section is structured to cover the datasets used, the evaluation metrics employed, and the downstream tasks for which the models were tested.
@@ 4.1 Datasets
Dataset # Ent # Rel # Train # Dev # Test Task Year
FIGER [44] - - 2,000,000 10,000 563 ET 2015
Open Entity [45] - - 2,000 2,000 2,000 ET 2018
TACRED [46] - 42 68,124 22,631 15,509 RC 2017
FewRel [47] - 80 8,000 16,000 16,000 RC 2018
WN11 [48] 38,696 11 112,581 2,609 10,544 TC 2013
FB13 [48] 75,043 13 316,232 5,908 23,733 TC 2013
WN9 [49] 6,555 9 11,741 1,337 1,319 TC 2022
FB15K-237N [50] 13,104 93 87,282 7,041 8,226 TC 2022
Nations [51] 14 55 1,592 199 201 LP 2007
FB15K [52] 14,951 1,345 483,142 50,000 59,071 LP 2013
FB15k-237 [53] 14,541 237 272,115 17,535 20,466 LP 2015
WN18RR [54] 40,943 11 86,835 3,034 3,134 LP, LP(ZS) 2018
UMLS [54] 135 46 5,216 652 661 LP 2018
YAGO3-10 [54] 103,222 30 490,214 2,295 2,292 LP 2018
NELL-ONE [55] 68,545 822 189,635 1,004 2,158 LP(ZS) 2018
CoDEx-S [56] 45,869 68 32,888 1,827 1,828 LP 2020
CoDEx-M [56] 11,941 50 185,584 10,310 10,311 LP 2020
CoDEx-L [56] 45,869 69 551,193 30,622 30,622 LP 2020
Wikidata5M-transductive [28] 4,594,485 822 20,614,279 5,163 5,133 LP 2021
Wikidata5M-inductive [28] 4,594,458 822 20,496,514 6,699 6,894 LP(ZS) 2021
Diabetes [57] 7,886 67 56,830 1,344 1,936 LP 2022
@@ 4.3.1 Entity Typing
Model Year FIGER Open Entity
Acc Macro Micro P R F1
NFGEC [58] 2016 55.6 75.15 71.73 68.8 53.3 60.1
UFET [45] 2018 - - - 77.4 60.6 68
BERT [1] 2018 52.04 75.16 71.63 76.37 70.96 73.56
RoBERTa [6] 2019 - - - 77.4 73.6 75.4
ERNIE [59] 2019 57.19 76.51 73.39 78.42 72.9 75.56
KnowBert [60] 2019 - - - 78.7 72.7 75.6
KEPLER [28] 2021 - - - 77.8 74.6 76.2
@@ 4.3.2 Relation Classification
Model Year FewRel TACRED
P R F1 P R F1
CNN [61] 2015 69.51 69.64 69.35 70.3 54.2 61.2
PA-LSTM [62] 2017 - - - 65.7 64.5 65.1
C-GCN [63] 2018 - - - 69.9 63.3 66.4
BERT_Base [1] 2018 85.05 85.11 84.89 67.23 64.81 66
BERT_Large [1] 2018 - - - - - 70.1
RoBERTa [6] 2019 - - - 70.4 71.1 70.7
MTB [64] 2019 - - - - - 71.5
ERNIE [59] 2019 88.49 88.44 88.32 69.97 66.08 67.97
KnowBert [60] 2019 - - - 71.6 71.4 71.5
KEPLER [28] 2021 - - - 71.5 72.5 72
@@ 4.3.3 Relation Prediction
Model Year FB15K YAGO3-10-100
MR Hits@1 Hits@1
TransE [52] 2013 2.5 84.3 -
TransR [65] 2015 2.1 91.6 -
DKRL [66] 2016 2 90.8 -
TKRL [67] 2016 1.7 92.8 -
PTransE [68] 2015 1.2 93.6 -
SSP [69] 2017 1.2 - -
ProjE [70] 2017 1.2 95.7 -
KG-BERT [19] 2019 1.2 96 -
KGT5 [34] 2022 - - 60
ChatGPT 2023 - - 39
GPT-4 2023 - - 56
LLaMA-7B [36] 2023 - - 13
LLaMA-13B [36] 2023 - - 1
KG-LLM [40] 2023 - - 71
@@ 4.3.4 Triple Classification
Model Year WN11 FB13 UMLS CoDEX-S FB15k-237N
Acc Acc Acc Acc Acc
NTN [48] 2013 86.2 90 - - -
TransE [52] 2013 75.9 81.5 78.1 72.1 -
DistMult [71] 2014 87.1 86.2 86.8 66.8 -
ComplEx [72] 2016 - - 90.8 67.8 65.7
RotatE [73] 2019 - - 92.1 75.7 68.5
KG-BERT [19] 2019 93.5 90.4 89.7 77.3 56
KG-GPT2 [43] 2021 85 89 - - -
LaSS [74] 2021 94.5 91.8 - - -
KGT5 [34] 2022 72.8 66.3 - - -
LMKE [75] 2022 - 91.7 92.4 - -
GPT-3.5 2023 - - 67.6 54.7 60.2
LLaMA-7B [36] 2023 21.1 9.1 - - -
LLaMA-13B [36] 2023 28.1 17.6 - - -
KG-LLM [40] 2023 95.6 90.2 86 80.3 80.5
KoPA [42] 2023 - - 92.6 82.7 77.7
@@ 4.3.5 Link Prediction
Table 9: Experimental statistics of WN18RR dataset under the LP subtask in transductive settings.
Model Year WN18RR
MR MRR Hits@1 Hits@3 Hits@10
TransE [52] 2013 2300 24.3 4.3 44.1 53.2
TransH [76] 2014 2524 - - - 50.3
DistMult [71] 2014 3704 44.4 41.2 47 50.4
TransR [65] 2015 3166 - - - 50.7
TransD [77] 2015 2768 - - - 50.7
ComplEx [72] 2016 3921 44.9 40.9 46.9 53
ConvE [54] 2018 4464 45.6 41.9 47 53.1
ConvKB [78] 2018 2554 24.9 - - 52.5
R-GCN [79] 2018 6700 12.3 8 13.7 20.7
KBGAN [80] 2018 - 21.5 - - 48.1
RotatE [73] 2019 3340 47.6 42.8 49.2 57.1
KBAT [81] 2019 1921 41.2 - - 55.4
CapsE [82] 2019 718 41.5 - - 55.9
QuatE [83] 2019 3472 48.1 43.6 50 56.4
TuckER [84] 2019 - 47 44.3 48.2 52.6
HAKE [85] 2019 - 49,7 45.2 51.6 58.2
AttH [86] 2020 - 48.6 44.3 49.9 57.3
REFE [87] 2020 - - - - 56.1
GAATs [88] 2020 1270 - - - 60.4
Complex-DURA [89] 2020 57.1 - - - -
DensE [90] 2020 3052 49.1 44.3 50.8 57.9
LineaRE [91] 2020 1644 49.5 45.3 50.9 57.8
RESCAL-DURA [92] 2020 - 49.8 45.5 - 57.7
CompGCN [93] 2020 - 47.9 44.3 49.4 54.6
NePTuNe [94] 2021 - - - - 55.7
ComplEx-N3-RP [95] 2021 - - - - 58
ConE [96] 2021 - 49.6 45.3 51.5 57.9
Rot-Pro [97] 2021 - 45.7 39.7 48.2 57.7
QuatDE [98] 2021 1977 48.9 43.8 50.9 58.6
NBFNet [99] 2021 - 55.1 49.7 - 66.6
KG-BERT [19] 2019 97 21.6 4.1 30.2 52.4
MTL-KGC [20] 2020 89 33.1 20.3 38.3 59.7
Pretrain-KGE [100] 2020 - 48.8 43.7 50.9 58.6
StAR [25] 2021 51 40.1 24.3 49.1 70.9
MEM-KGC [101] 2021 - 57.2 48.9 62 72.3
LaSS [74] 2021 35 - - - 78.6
SimKGC [26] 2021 - 66.7 58.8 72.1 80.5
LP-BERT [27] 2022 92 48.2 34.3 56.3 75.2
KGT5 [34] 2022 - 54.2 50.7 - 60.7
OpenWorld KGC [102] 2022 - 55.7 47.5 60.4 70.4
LMKE [75] 2022 79 61.9 52.3 67.1 78.9
GenKGC [32] 2022 - - 28.7 40.3 53.5
KG-S2S [35] 2022 - 57.4 53.1 59.5 66.1
kNN-KGE [103] 2023 - 57.9 52.5 - -
CSPromp-KG [24] 2023 - 57.5 52.2 59.6 67.8
GPT-3.5 2023 - - 19 - -
LLaMA-7B [36] 2023 - - 8.5 - -
LLaMA-13B [36] 2023 - - 10 - -
KG-LLM [40] 2023 - - 25.59 - -
CD [38] 2024 - 57.6 52.6 60.7 67.2
CP-KGC [39] 2024 - 67.3 59.9 72.1 80.4
KICGPT [41] 2024 - 56.4 47.8 61.2 67.7